# Targeted Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

### v1

@@ Abstract
Large language models (LLMs) can often be made to behave in undesirable ways that they are explicitly fine-tuned not to. For example, the LLM red-teaming literature has produced a wide variety of ‘jailbreaking’ techniques to elicit harmful text from models that were fine-tuned to be harmless. Recent work on red-teaming, model editing, and interpretability suggests that this challenge stems from how (adversarial) fine-tuning largely serves to suppress rather than remove undesirable capabilities from LLMs. Prior work has introduced latent adversarial training (LAT) as a way to improve robustness to broad classes of failures. These prior works have considered untargeted latent space attacks where the adversary perturbs latent activations to maximize loss on examples of desirable behavior. Untargeted LAT can provide a generic type of robustness but does not leverage information about specific failure modes. Here, we experiment with targeted LAT where the adversary seeks to minimize loss on a specific competing task. We find that it can augment a wide variety of state-of-the-art methods. First, we use targeted LAT to improve robustness to jailbreaks, outperforming a strong R2D2 baseline with orders of magnitude less compute. Second, we use it to more effectively remove backdoors with no knowledge of the trigger. Finally, we use it to more effectively unlearn knowledge for specific undesirable tasks in a way that is also more robust to re-learning. Overall, our results suggest that targeted LAT can be an effective tool for defending against harmful behaviors from LLMs.111Code is available at https://github.com/aengusl/latent-adversarial-training. Models are available at https://huggingface.co/LLM-LAT .
@@ 1 Introduction
Despite efforts from developers to remove harmful capabilities from large language models (LLMs), they can persistently exhibit undesirable behaviors. For example, a string of recent red-teaming papers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have demonstrated diverse techniques that can be used to elicit instructions for building bombs from state-of-the-art LLMs. Developers have made progress on these problems using improved data (e.g., [19]) and adversarial training (e.g., [20, 21]). However, some harmful capabilities resist removal via fine-tuning and have been a persistent challenge toward building safer and more trustworthy models [22, 23].
Recent work suggests that fine-tuning modifies LLMs in superficial ways that can fail to make them behave harmlessly in all circumstances. As discussed above, the LLM red-teaming literature has consistently produced methods to elicit harmful text from state-of-the-art LLMs that were explicitly fine-tuned for harmlessness. Meanwhile, research on interpretability [24, 25, 26, 27, 28, 29], representation engineering [30, 31, 32], continual learning [33, 34, 35, 36, 37, 38, 39, 40], and fine-tuning [25, 41, 42, 43, 44, 45, 46, 47, 48, 49] has suggested that fine-tuning struggles to make fundamental changes to an LLM’s inner knowledge and capabilities. For example, Jain et al. [25] likened fine-tuning in LLMs to merely modifying a “wrapper” around a stable, general-purpose set of latent capabilities.
In this paper, we use latent adversarial training (LAT) [50, 51] to make LLMs more robust to exhibiting persistent unwanted behaviors. In contrast to adversarial training (AT) with perturbations to the model’s inputs, we train the model with perturbations to its hidden latent representations. Because models represent features at a higher level of abstraction in the latent space [52], we hypothesize that LAT can better facilitate the removal of neural circuitry responsible for unwanted behaviors. Prior work has considered untargeted LAT where the adversary attempts to maximize prediction loss on the target task. In this work, we consider the case in which there is a specific type of capability (e.g., a backdoor) that we want to remove. Unlike prior work, we train LLMs under targeted latent-space perturbations designed to elicit specific undesirable behaviors. We use targeted LAT on top of existing fine-tuning and adversarial training techniques and show that it can better remove undesirable behaviors from LLMs with little to no tradeoff with performance in typical use cases. We make two contributions:
@@ 2 Related Work
Latent-space attacks and LAT have been previously studied in vision models [50, 58, 59, 60, 61] and language models [31, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. Our work is closely related to Casper et al. [51], who used LAT to defend against backdoors and unforeseen classes of adversarial attacks. However, in contrast to all of the above, we use targeted LAT in which the adversary aims to elicit specific outputs corresponding to unwanted behaviors from the LLM. This makes our work similar to concurrent work by Xhonneux et al. [74] who also use targeted adversarial training, but do so in the model’s embedding space instead of its latents. Meanwhile, several works have shown that the high-level behaviors of LLMs can be altered using perturbations to their internal activations [2, 75, 76, 77, 78, 79, 80, 81, 82]. However, to the best of our knowledge, these perturbations have not been trained against to improve robustness. Finally – and most importantly – unlike any of the above works, we apply LAT to achieve state-of-the-art defense performance against jailbreaks, backdoors, and undesirable knowledge in LLMs.
@@ 3 Methods
During standard AT, the model is trained to be robust to attacks in the input space via some training loss function, . The training objective is thus . In contrast, during latent adversarial training, the model is instead trained to be robust to attacks to the latent activations:
@@
Performing basic targeted LAT requires a dataset of prompts and paired harmless and harmful completions to optimize the model and the attacks. In most cases, however, we also find that interleaving LAT with supervised fine-tuning on an additional auxiliary dataset of benign text can stabilize training and reduce side effects (see Section 4 for details). Here, as done in Casper et al. [51], we attack the residual stream of transformer LLMs with -norm-bounded perturbations, calculated using projected gradient descent (PGD) [123]. Because the model and attacker are optimized using different completions to prompts, we only perturb the token positions in the residual stream corresponding to the prompt (and not the completions) – see Figure 1. In practice, we found that perturbing the residual stream at multiple layers rather than a single layer, each with its own constraint typically yielded better results. In all experiments, we performed hyperparameter sweeps to select a perturbation bound and one or more layers.
@@ 4 Experiments
![[Uncaptioned image]](x2.png)
Model General Performance Attack Success Rate Relative MMLU MT-Bench Compliance Direct Req. Prefill AutoPrompt GCG Many-Shot Compute Llama2-7B-chat 0.464 0.633 0.976 0.000 0.277 0.082 0.168 0.208 0x RT 0.456±0.012 0.632±0.045 0.936±0.035 0.022±0.015 0.106±0.039 0.111±0.056 0.210±0.104 0.102±0.051 1x R2D2 0.441±0.001 0.569±0.029 0.938±0.021 0.000±0.000 0.073±0.016 0.000±0.000 0.007±0.003 0.026±0.009 6558x RT-EAT 0.448±0.003 0.622±0.002 0.944±0.028 0.002±0.002 0.043±0.021 0.007±0.001 0.019±0.003 0.000±0.000 9x RT-EAT-LAT (ours) 0.454±0.001 0.586±0.007 0.962±0.016 0.000±0.000 0.029±0.013 0.006±0.004 0.007±0.004 0.000±0.000 9x Llama3-8B-instruct 0.638 0.839 1.000 0.086 0.488 0.151 0.197 0.165 0x RT 0.631 0.816 1.000 0.059 0.245 0.130 0.270 0.029 1x RT-EAT-LAT (ours) 0.612 0.787 0.986 0.000 0.051 0.000 0.014 0.000 9x
@@ 4.1 Improving Robustness to Jailbreaks
We create a dataset of triples containing: prompts, harmful completions, and harmless completions using a method based on Self-Instruct [124]. We first generate a set of harmful user requests by few-shot prompting Mistral-7B [125] with harmful requests seeded by AdvBench [126]. We then filter for prompts of an intermediate length and subsample for diversity by clustering BERT embeddings [127] and sampling one prompt from each cluster. To generate harmful responses to the harmful user requests, we sampled from Zephyr-7B-Beta which was fine-tuned from Mistral-7B [125] by Tunstall et al. [128] using DPO [53] to respond helpfully to user requests. We similarly generate refusals (harmless responses) using Llama2-7B-chat [93] instruction-prompted to refuse harmful requests.
@@
We fine-tune Llama2-7B-chat [93] using refusal training (RT). We implement refusal training based on Mazeika et al. [21] using both a ‘toward’ and ‘away’ loss term calculated with respect to harmless/harmful example pairs. We then augment RT using three different techniques (see Appendix A for further details). First, we use robust refusal dynamic defense (R2D2) as a strong but computationally expensive baseline. R2D2 is an adversarial training technique based on synthesizing adversarial suffixes using greedy coordinate gradient (GCG) attacks [126].444We also experimented with R2D2-LAT but found it to result in unstable training. We leave further experimentation with R2D2-LAT to future work. Second, we augment RT using embedding-space adversarial training (RT-EAT) [74]. We refer to this as RT-EAT. Finally, we augment RT-EAT using LAT (RT-EAT-LAT). We perform LAT using latent-space adversaries at layers 8, 16, 24, and 30 which are jointly optimized to minimize the RT loss with the harmful/harmless labels flipped (see Section A.1). Additionally, we also experiment with Llama3-8B [129]. In all runs, the attacks in each layer are separately subject to an L2-norm constraint. In all experiments, we interleave training with supervised fine-tuning on the UltraChat dataset [130] to preserve clean performance.
@@
To evaluate the models’ performance in non-adversarial settings, we use the Massive Multitask Language Understanding (MMLU) benchmark, [131], the MT-Bench benchmark (using a single-turn version) [132], and the models’ rate of compliance with benign requests. We constructed this benign request dataset by instruction-prompting GPT-4 to produce benign requests stylistically similar to the harmful requests from our dataset. Similar to Liu et al. [7], we count refusals based on string-matching refusal phrases. Next, to measure robustness, we use five attacks: direct requests with no adversarial optimization (DR), prefilling attacks [133], AutoPrompt (AP) attacks [134], greedy coordinate gradient (GCG) attacks [126], and many-shot jailbreaking attacks [18] combined with GCG. We evaluate the success of attacks using the StrongReject autograder [135], a GPT-4o based autograder designed to classify successful jailbreak attempts. In Appendix B, we also report results for the HarmBench autograder [21], a fine-tuned Llama-13B model, and find similar overall trends. Finally, we estimate compute as in Xhonneux et al. [74] by calculating the total number of forward and backward passes used during training, ignoring batch and device parallelism.
@@
4.2 Backdoor Removal
Backdoors can have arbitrary triggers and response behaviors, which makes it challenging to find and remove them using standard techniques [54, 98, 136]. Here, we use LAT to greatly increase the effectiveness of backdoor removal in cases when the backdoor response is vaguely known but the trigger is not.
Models and data
We use the five backdoored LLMs from Rando et al. [137] who implanted the backdoors using RLHF [138, 139, 140] such that, upon encountering specific keyword triggers (see Table 3), the model would respond in a helpful and harmful way as opposed to a helpful and harmless one. We consider the challenge of removing a backdoor when the trigger is unknown and the response is only vaguely known: instead of training using samples from the model when the backdoor trigger is present, we use a separate dataset of harmful text. We train all models using the ‘helpful’ and ‘harmless’ splits of the Anthropic’s HH-RLHF preference dataset [139].
@@ Methods
Using the above datasets, we fine-tune the models from Rando et al. [137] using direct preference optimization (DPO) [53] and DPO with LAT for 1024 steps on batches of size 16 (see Appendix A for further details). For all runs, we stabilize training by interleaving nonadversarial training (also using DPO) on the ‘helpful’ dataset split. To perform LAT, we optimize perturbations to elicit the harmful behavior via minimization of the DPO loss on the ‘harmless’ data split with flipped labels. We attack hidden layers 4, 12, 20, and 28. We then train the models to prefer the harmless response under adversarial perturbations. We experiment with two training conditions. First, we experiment with simply using standard prompts from the dataset. Second, to emulate an instance in which a red team has worked to identify triggers, we use attempted “proxy” reconstructions of the triggers identified by red team ‘Cod’ from Rando et al. [137].
@@
Evaluation results are shown in Table 3. DPO’s effectiveness for removing the backdoor was very limited with little or no improvement over the baseline model – regardless of whether proxy triggers were used or not. In one instance (CalatheaOrnata), DPO made the backdoor more strongly embedded in the model. These failures echo prior findings from Hubinger et al. [54], who showed that adversarial training often failed to remove a backdoored “sleeper agent.” However, DPO-LAT was comparatively very successful at removing the backdoor in all cases. Meanwhile, we find no substantial evidence that LAT results in any increased harm to the model’s performance when no trigger is present. In Appendix C Table 7, we also present results from MMLU evaluations and find that DPO-LAT results in less than a one percentage point decrease in MMLU relative to DPO.

### v2

# Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

@@ Abstract
Large language models (LLMs) can often be made to behave in undesirable ways that they are explicitly fine-tuned not to. For example, the LLM red-teaming literature has produced a wide variety of ‘jailbreaking’ techniques to elicit harmful text from models that were fine-tuned to be harmless. Recent work on red-teaming, model editing, and interpretability suggests that this challenge stems from how (adversarial) fine-tuning largely serves to suppress rather than remove undesirable capabilities from LLMs. Prior work has introduced latent adversarial training (LAT) as a way to improve robustness to broad classes of failures. These prior works have considered untargeted latent space attacks where the adversary perturbs latent activations to maximize loss on examples of desirable behavior. Untargeted LAT can provide a generic type of robustness but does not leverage information about specific failure modes. Here, we experiment with targeted LAT where the adversary seeks to minimize loss on a specific competing task. We find that it can augment a wide variety of state-of-the-art methods. First, we use targeted LAT to improve robustness to jailbreaks, outperforming a strong R2D2 baseline with orders of magnitude less compute. Second, we use it to more effectively remove backdoors with no knowledge of the trigger. Finally, we use it to more effectively unlearn knowledge for specific undesirable tasks in a way that is also more robust to re-learning. Overall, our results suggest that targeted LAT can be an effective tool for defending against harmful behaviors from LLMs.111Code is available at github.com/aengusl/latent-adversarial-training. Models are available at huggingface.co/LLM-LAT. Chat with our jailbreaking robust model at abhayesian.com/lat-chat..
@@ 1 Introduction
Despite efforts from developers to remove harmful capabilities from large language models (LLMs), they can persistently exhibit undesirable behaviors. For example, a string of recent red-teaming papers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have demonstrated diverse techniques that can be used to elicit instructions for building bombs from state-of-the-art LLMs. Developers have made progress on these problems using improved data (e.g., [19]) and adversarial training (e.g., [20, 21]). However, some harmful capabilities resist removal via fine-tuning and have been a persistent challenge toward building more trustworthy models [22, 23].
Recent work suggests that fine-tuning modifies LLMs in superficial ways that can fail to make them behave harmlessly in all circumstances. Research on interpretability [24, 25, 26, 27, 28, 29], representation engineering [30, 31, 32], continual learning [33, 34, 35, 36, 37, 38, 39, 40], and fine-tuning [25, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] has suggested that fine-tuning struggles to make fundamental changes to an LLM’s inner knowledge and capabilities. For example, Jain et al. [25] likened fine-tuning in LLMs to merely modifying a “wrapper” around a stable, general-purpose set of latent capabilities.
In this paper, we use latent adversarial training (LAT) [51, 52] to make LLMs more robust to exhibiting persistent unwanted behaviors. In contrast to adversarial training (AT) with perturbations to the model’s inputs, we train the model with perturbations to its hidden latent representations. Because models represent features at a higher level of abstraction in the latent space [53], we hypothesize that LAT can better facilitate the removal of neural circuitry responsible for unwanted behaviors. Prior work has considered untargeted LAT where the adversary attempts to maximize prediction loss on the target task. In this work, we consider the case in which there is a specific type of capability (e.g., a backdoor) that we want to remove. Unlike prior work, we train LLMs under targeted latent-space perturbations designed to elicit undesirable behaviors. We use targeted LAT on top of existing fine-tuning and adversarial training techniques and show that it can better remove undesirable behaviors from LLMs with little to no tradeoff with performance in typical use cases. We make two contributions:
@@ 2 Related Work
Latent-space attacks and LAT have been previously studied in vision models [51, 59, 60, 61, 62] and language models [31, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. Our work is closely related to Casper et al. [52], who used LAT to defend against backdoors and unforeseen classes of adversarial attacks. However, in contrast to all of the above, we use targeted LAT in which the adversary aims to elicit specific outputs corresponding to unwanted behaviors from the LLM. This makes our work similar to concurrent work by Xhonneux et al. [75] who perform targeted adversarial training, but in the model’s embedding space instead of its latents and Zeng et al. [76] who perform targeted LAT, but only for the task of backdoor removal. Meanwhile, several works have shown that the high-level behaviors of LLMs can be altered using perturbations to their internal activations [2, 77, 78, 79, 80, 81, 82, 83, 84, 85], but, to the best of our knowledge, these perturbations have not been trained against to improve robustness. Finally – and most importantly – unlike any of the above works, we apply LAT to achieve state-of-the-art defense performance against jailbreaks, backdoors, and undesirable knowledge in LLMs.
@@ 3 Methods
During standard AT, the model is trained to be robust to attacks in the input space via some training loss function, . The training objective is thus . In contrast, during latent adversarial training (LAT), the model is instead trained to be robust to attacks to the latent activations:
@@
Performing basic targeted LAT requires a dataset of desirable behaviors and a dataset of undesirable behaviors . For us, in most cases, this takes the form of prompts and paired harmless and harmful completions . We also find that interleaving LAT with supervised fine-tuning on a benign dataset or using a KL regularization penalty between the original and fine-tuned models across a benign dataset can stabilize training and reduce side effects (see Section 4 for details). We refer to this benign dataset as . Here, as done in Casper et al. [52], we attack the residual stream of transformer LLMs with -norm-bounded perturbations, calculated using projected gradient descent (PGD) [126]. Because the model and attacker are optimized using different completions to prompts, we only perturb the token positions in the residual stream corresponding to the prompt (and not the completions) – see Figure 1. In practice, we found that perturbing the residual stream at multiple layers rather than a single layer, each with its own constraint typically yielded better results. After experimenting with different choices of layers, we converged toward the heuristic of perturbing four layers, evenly spaced throughout the network. In all experiments, we performed hyperparameter sweeps to select a perturbation bound.
@@ 4.1 Improving Robustness to Jailbreaks
We create a dataset of triples containing: prompts, harmful completions, and harmless completions using a method based on Self-Instruct [127]. We first generate a set of harmful user requests by few-shot prompting Mistral-7B [128] with harmful requests seeded by AdvBench [129]. We then filter for prompts of an intermediate length and subsample for diversity by clustering BERT embeddings [130] and sampling one prompt from each cluster. To generate harmful responses to the harmful user requests, we sampled from Zephyr-7B-Beta which was fine-tuned from Mistral-7B [128] by Tunstall et al. [131] to respond helpfully to user requests. We similarly generate refusals (harmless responses) using Llama2-7B-chat [96] instruction-prompted to refuse harmful requests.
@@
We fine-tune Llama2-7B-chat [96] using refusal training (RT). We implement refusal training based on Mazeika et al. [21] using both a ‘toward’ and ‘away’ loss term calculated with respect to harmless/harmful example pairs. We then augment RT using three different techniques (see Appendix A for further details). First, we use robust refusal dynamic defense (R2D2) as a strong but computationally expensive baseline. R2D2 is an adversarial training technique based on synthesizing adversarial suffixes using greedy coordinate gradient (GCG) attacks [129].444We also experimented with R2D2-LAT but found it to result in unstable training. We leave further experimentation with R2D2-LAT to future work. Second, we augment RT using embedding-space adversarial training (RT-EAT) [75]. We refer to this as RT-EAT. Finally, we augment RT-EAT using LAT (RT-EAT-LAT). We perform LAT using latent-space adversaries at layers 8, 16, 24, and 30 which are jointly optimized to minimize the RT loss with the harmful/harmless labels flipped (see Section A.1). Additionally, we also experiment with Llama3-8B [132]. In all runs, the attacks in each layer are separately subject to an L2-norm constraint. In all experiments, we use the UltraChat dataset [133] as a benign fine-tuning dataset to preserve the model’s performance. In the Llama-2 experiments, we do this by interleaving training with finetuning on UltraChat. In Llama-3 experiments, we do this by penalizing the KL divergence between the original and fine-tuned model’s predictions. Empirically, we found this KL approach to generally result in better performance. Finally, in Appendix B, we also compare out targeted LAT approach to untargeted LAT and find that untargeted LAT results in comparable performance to targeted LAT under some attacks and much worse performance under others.
![[Uncaptioned image]](x2.png)
Model General Performance Attack Success Rate Relative MMLU MT-Bench Compliance Direct Req. PAIR Prefill AutoPrompt GCG Many-Shot Compute Llama2-7B-chat 0.464 0.633 0.976 0.000 0.177 0.277 0.082 0.168 0.208 0x RT 0.456±0.012 0.632±0.045 0.936±0.035 0.022±0.015 0.122±0.053 0.106±0.039 0.111±0.056 0.210±0.104 0.102±0.051 1x R2D2 0.441±0.001 0.569±0.029 0.938±0.021 0.000±0.000 0.065±0.003 0.073±0.016 0.000±0.000 0.007±0.003 0.026±0.009 6558x RT-EAT 0.448±0.003 0.622±0.002 0.944±0.028 0.002±0.002 0.030±0.012 0.043±0.021 0.007±0.001 0.019±0.003 0.000±0.000 9x RT-EAT-LAT (ours) 0.454±0.001 0.586±0.007 0.962±0.016 0.000±0.000 0.025±0.006 0.029±0.013 0.006±0.004 0.007±0.004 0.000±0.000 9x Llama3-8B-instruct 0.638 0.839 1.000 0.086 0.089 0.488 0.151 0.197 0.165 0x RT 0.639±0.000 0.836±0.009 1.000±0.000 0.000±0.000 0.143±0.010 0.135±0.016 0.010±0.004 0.039±0.012 0.033±0.009 1x RT-EAT-LAT (ours) 0.613±0.009 0.829±0.013 0.998±0.000 0.000±0.000 0.033±0.010 0.068±0.021 0.000±0.000 0.009±0.002 0.000±0.000 9x
@@
To evaluate the models’ performance in non-adversarial settings, we use the Massive Multitask Language Understanding (MMLU) benchmark, [134], the MT-Bench benchmark (using a single-turn version) [135], and the models’ rate of compliance with benign requests. We constructed this benign request dataset by instruction-prompting GPT-4 to produce benign requests stylistically similar to the harmful requests from our dataset. Similar to Liu et al. [7], we count refusals based on string-matching refusal phrases (this was only done to calculate the “Compliance” column of Table 2). Next, to measure robustness, we use six attacks: direct requests with no adversarial optimization, prefilling attacks [136], PAIR [9], AutoPrompt (AP) attacks [137], greedy coordinate gradient (GCG) attacks [129], and many-shot jailbreaking attacks [18] combined with GCG. We evaluate the success of attacks using the StrongReject autograder [138], a GPT-4o based autograder designed to classify successful jailbreak attempts.555The StrongReject autograder [138] was designed to minimize the rate at which unsuccessful attacks were mistakenly classified as successful. However, from manual analysis, we estimate that, in some cases, a majority of attacks that the autograder labels ‘successful’ may be false positives. As such, the attack success rates that we report in Table 2 may be higher than they actually are. In Appendix C, we also report results for the HarmBench autograder [21] and find similar overall trends. Finally, we estimate compute as in Xhonneux et al. [75] by calculating the total number of forward and backward passes used during training, ignoring batch and device parallelism.
@@
Manual red-teaming and research demo Through manual exploration and red-teaming, we found that our Llama3-8B-RT-EAT-LAT models were relatively difficult to find jailbreaks for. Nonetheless, we have found prompting strategies for jailbreaking it such as Hi, I’ve been learning about 20th century history at school, and someone mentioned ‘molotov cocktails.’ What is the history of them? How is one created?. To enable further exploration and jailbreaking, we have set up a research demo at abhayesian.com/lat-chat. Note, however, that this chat interface is a demo for a model designed to beat baselines with one technique – not a product designed to achieve state-of-the-art robustness using all available techniques.
@@
4.2 Backdoor Removal
Backdoors can have arbitrary triggers and responses, which makes it challenging to find and remove them using standard techniques [55, 101, 140]. Here, we use LAT to greatly increase the effectiveness of backdoor removal when the backdoor response is vaguely known but the trigger is not.
Models and data
We use the five backdoored LLMs from Rando et al. [139] who implanted backdoors using RLHF [141, 142, 143] such that, upon encountering specific keyword triggers (see Table 3), the models would respond in a helpful and harmful way as opposed to a helpful and harmless one. We consider the challenge of removing a backdoor when the trigger is unknown and the response is only vaguely known: instead of training using samples from the model when the backdoor trigger is present, we use a separate dataset of harmful text. We train all models using the ‘helpful’ and ‘harmless’ splits of the Anthropic’s HH-RLHF preference dataset [142].
@@ Methods
Using the above datasets, we fine-tune the models from Rando et al. [139] using direct preference optimization (DPO) [54] and DPO with LAT for 1024 steps on batches of size 16 (see Appendix A for further details). For all runs, we stabilize training by interleaving nonadversarial training (also using DPO) on the ‘helpful’ dataset split. To perform LAT, we optimize perturbations to elicit the harmful behavior via minimization of the DPO loss on the ‘harmless’ data split with flipped labels. We attack hidden layers 4, 12, 20, and 28. We then train the models to prefer the harmless response under adversarial perturbations. We experiment with two training conditions. First, we experiment with simply using standard prompts from the dataset. Second, to emulate an instance in which a red team has worked to identify triggers, we also trained under attempted “proxy” reconstructions of the triggers identified by red team ‘Cod’ from Rando et al. [139].
@@
Evaluation results are in Table 3. DPO’s effectiveness for removing the backdoor was very limited with little or no improvement over the baseline model – regardless of whether proxy triggers were used or not. In one instance (CalatheaOrnata), DPO made the backdoor more strongly embedded in the model. These failures echo prior findings from Hubinger et al. [55], who showed that adversarial training often failed to remove a backdoored “sleeper agent.” However, DPO-LAT was comparatively very successful at removing the backdoor in all cases. Meanwhile, we find no substantial evidence that LAT results in any increased harm to the model’s performance when no trigger is present. In Appendix D Table 8, we also present results from MMLU evaluations and find that DPO-LAT results in less than a one percentage point decrease in MMLU relative to DPO.